Which has more number of atoms, 100 grams of sodium or 100 grams of iron (given atomic mass of Na = 23 u, Fe = 56 u)?

Grade:12

2 Answers

Harshit Singh
3 years ago
Dear student


23 g of Na = 6.022 × 1023 atoms (1 mole).
100 g of Na = ?
100 x 6.022 x 10^23​
56
602.2x 10^23
23

= 26.182 ×10^23
= 2.6182×10^24atoms
56 g of Fe = 6.022 × 10^23atoms
100 g o Fe = ?
100 g of Na contain → 2.618 × 10^24atoms
100 g of Fe contain → 1.075 × 10^24atoms
∴ 100 g of Na contains more atoms.



Thanks
Harshit Singh
3 years ago
Dear student


23 g of Na = 6.022 × 1023 atoms (1 mole).
100 g of Na = ?
100 x 6.022 x 10^23​
56
602.2x 10^23
23

= 26.182 ×10^23
= 2.6182×10^24atoms
56 g of Fe = 6.022 × 10^23atoms
100 g o Fe = ?
100 g of Na contain → 2.618 × 10^24atoms
100 g of Fe contain → 1.075 × 10^24atoms
∴ 100 g of Na contains more atoms.



Thanks